At the beginning of the year, Palermo Brothers, Inc., purchased a new plastic water bottle making machine at a cost of $78,000. The estimated residual value was $9,000. Assume that the estimated useful life was four years, and the estimated productive life of the machine was 690,000 units. Actual annual production was as follows:

Year Units
1 207,000
2 151,800
3 189,750
4 141,450


Required:

1. Complete a separate depreciation schedule for each of the alternative methods.


a. Straight-line.

b. Units-of-production.

c. Double-declining-balance.
Answer #1

Answers

  • Requirement [a] Straight Line Method

Working

A

Cost

$            78,000.00

B

Residual Value

$              9,000.00

C=A - B

Depreciable base

$            69,000.00

D

Life [in years]

4

E=C/D

Annual SLM depreciation

$            17,250.00

Depreciation Schedule

Year

Book Value

Depreciation expense

Ending Book Value

Accumulated Depreciation

1

$              78,000.00

$            17,250.00

$         60,750.00

$             17,250.00

2

$              60,750.00

$            17,250.00

$         43,500.00

$             34,500.00

3

$              43,500.00

$            17,250.00

$         26,250.00

$             51,750.00

4

$              26,250.00

$            17,250.00

$            9,000.00

$             69,000.00

  • Requirement [b] Units of production

Working

A

Cost

$            78,000.00

B

Residual Value

$              9,000.00

C=A - B

Depreciable base

$            69,000.00

D

Usage

                   690,000

E = C/D

Depreciation per unit

$                       0.10

Depreciation Schedule

Year

Book Value

Usage

Depreciation expense

Ending Book Value

Accumulated Depreciation

1

$              78,000.00

                   207,000

$         20,700.00

$             57,300.00

$        20,700.00

2

$              57,300.00

                   151,800

$         15,180.00

$             42,120.00

$        35,880.00

3

$              42,120.00

                   189,750

$         18,975.00

$             23,145.00

$        54,855.00

4

$              23,145.00

                   141,450

$         14,145.00

$                9,000.00

$        69,000.00

  • Requirement [c] Double Declining balance

Working

A

Cost

$            78,000.00

B

Residual Value

$              9,000.00

C=A - B

Depreciable base

$            69,000.00

D

Life [in years]

4

E=C/D

Annual SLM depreciation

$            17,250.00

F=E/C

SLM Rate

25.00%

G=F x 2

DDB Rate

50.00%

Depreciation Schedule

Year

Beginning Book Value

Depreciation rate

Depreciation expense

Ending Book Value

Accumulated Depreciation

1

$              78,000.00

50.00%

$         39,000.00

$             39,000.00

$        39,000.00

2

$              39,000.00

50.00%

$         19,500.00

$             19,500.00

$        58,500.00

3

$              19,500.00

50.00%

$            9,750.00

$                9,750.00

$        68,250.00

4

$                9,750.00

$            750.00

$                9,000.00

$        69,000.00